If x is normally distributed with µ = 283 and σ = 21 ,find P(x > 360).
Answer #1

X is normally distributed with mean µ = 283 and σ = 21

Now we know that \frac{x-\mu}{\sigma} ~ N(0,1)

P(X > 360) = 1- P(x\leq 360)

= 1- P(\frac{x-\mu}{\sigma}\leq \frac{360-283}{21})

= 1- P(Z\leq 3.6667)

= 1- 0.99987715

= 0.00012285

P(X > 360) = 0.00012285